slove
log6 4x-log6 (x-3)=log6 12

Answer by stanbon(29620)

log6 4x-log6 (x-3)=log6 12
Keep in mind the base is 6:
=log(4x/(x-3)) = log12
Take the anti-log of both sides to get:
(4x/(x-3)) = 12
4x=12(x-3)
4x=12x-36
8x=36
x=9/2
Notice that the base 6 made no difference.
Cheers,
Stan H.